A car travels for 2.00 h with a constant velocity of 40.0 km/h East, and then for 3.00 h with a constant velocity of 70.0 km/h also East. Find the average velocity for the entire journey if the Eastern direction is taken as positive


1
Expert's answer
2021-10-01T12:22:11-0400

The average velocity is the total displacement over total time:


"<v>=\frac DT=\frac{v_1t_1+v_2t_2}{t_1+t_2}=58\text{ km/h}."